# Enriching Diagrams with Algebraic Operations

### 08 Dec 2023 v2

@@ 4 Enriched monoidal categories
A lax monoidal strict $2$-functor $(G,G^{0},G^{2})\colon(\mathbf{V},\otimes,I)\to(\mathbf{W},\times,\ast)$ induces a $2$-functor $\mathbf{PMon}(G)\colon\mathbf{PMon}(\mathbf{V},\otimes,I)\to\mathbf{PMon}(\mathbf{W},\times,\ast)$ between $2$-categories of pseudo-monoids, lax homomorphisms and $2$-cells that are compatible with the homomorphism structures. If $G$ has a monoidal left adjoint $F$, then $\mathbf{PMon}(F)$ is left adjoint to $\mathbf{PMon}(G)$. Finally, if the monoidal categories and functors are symmetric, then the adjunction can be improved to one between symmetric pseudo-monoids.
@@
From this theorem we can derive the following corollary, which is our main tool for building monoidal diagrams that are enriched with algebraic operations.
@@
For any monoidal affine monad $T$ on $\mathbf{Set}$, the free-underlying adjunction $(-)_{0}:\mathbf{Alg}^{T}\text{-}\mathbf{Cat}\rightleftarrows\mathbf{Cat}:F_{*}$ lifts to an adjunction between the $2$-categories of symmetric monoidal $\mathbf{Alg}^{T}$-enriched and $\mathbf{Set}$-enriched categories $(-)_{0}:\mathbf{Alg}^{T}\text{-}\mathbf{SMCat}\rightleftarrows\mathbf{SMCat}:F_{*}$. More explicitly, given a monoidal affine monad $T$ on $\mathbf{Set}$ and a SMC $(\mathbf{C},\otimes,I,\alpha,\lambda,\rho,\sigma)$ we can construct the freely enriched symmetric monoidal $\mathbf{Alg}^{T}$-category $(\underline{\mathbf{C}},\underline{\otimes},\underline{I},\underline{\alpha},\underline{\lambda},\underline{\rho},\underline{\sigma})$.
@@ 5.3 Enriching the categories $\mathbf{Qubit}$ and $\mathbf{CPM(\mathbf{Qubit})}$
We also define the non-freely enriched category $\underline{\mathbf{CPM(\mathbf{Qubit})}}$ so we can interpret probability distributions as CP maps. For this, we define SMC-structure in the non-free $(\mathbf{Alg}^{\mathcal{D}},\otimes_{\mathcal{D}},\mathcal{D}(\ast))$, with a tensor product of algebras defined by the coequalizer (1), with a more detailed description in Appendix 0.B. The category $(\underline{\mathbf{CPM(\mathbf{Qubit})}},\odot,U)$ has the same objects as $\mathbf{CPM(\mathbf{Qubit})}$ and for every pair of objects $A,B$ the hom-object is an algebra $(\mathbf{CPM(\mathbf{Qubit})}(A,B),\alpha)$. Composition of hom-objects follows from composition in $\mathbf{CPM(\mathbf{Qubit})}$, and for an object $A$ the identity element is $j_{A}:(\ast,\alpha)\to(\operatorname{Id}_{A},\alpha)$. We define now its symmetric monoidal structure following the definition of enriched SMC from the beginning of Section 4. The tensor product $\odot$ on objects is the same as in $\mathbf{CPM(\mathbf{Qubit})}$, and on hom-objects it is the tensor product in $\mathbf{CPM(\mathbf{Qubit})}$ to the underlying sets: $\odot:(\mathbf{CPM(\mathbf{Qubit})}(A,A^{\prime}),\alpha)\otimes_{\mathcal{D}}(\mathbf{CPM(\mathbf{Qubit})}(B,B^{\prime}),\alpha)$to $(\mathbf{CPM(\mathbf{Qubit})}(A\otimes B,A^{\prime}\otimes B^{\prime}),\alpha)$. The unit $U$ is the one in $\mathbf{CPM(\mathbf{Qubit})}$. The associator, unitors, and symmetry all follow from applying the ones in $\mathbf{CPM(\mathbf{Qubit})}$.
@@ 5.4 Enriched ZX-diagrams and their interpretation
We need to take extra care when handling scalars inside the brackets. Indeed, what we have inside distribution brackets is a formal convex sum of ZX-diagrams (or, in the general case, string diagrams), meaning that the SMC rewriting axioms apply to each summand independently. Since summands are also juxtaposed, it might seem like this notation allows for the transfer of scalars from one summand to another. The crux is that, since what is enclosed by trapezoids is a formal sum, we cannot drag scalars from one summand to another using those same monoidal category axioms. This means that we can consider the probabilities (and any scalar factor if present, such as the imaginary unit in Figure 5) to be bound to the wires themselves, and only interact with the ZX-diagrams (or generally string diagrams) that belong to that summand. An alternative done in [46, 36] is to encapsulate each summand in “bubbles”, which might help separating them visually more.
@@ 6 Discussion and future work
In this work, we have shown how to construct freely enriched symmetric monoidal categories over the algebras of a commutative affine monad. We have then taken this construction and developed a graphical language that captures the additional algebraic structure of the morphisms for the case of the Distribution monad. We then show how we can use this to study classical probabilistic processes in quantum systems, a highly relevant type of operation for near-term quantum applications. In particular, we extend the ZX-calculus to make it a language for reasoning in an enriched version of $\mathbf{CPM(\mathbf{Qubit})}$.
We believe that this work opens several directions for future research. The most evident one is to prove completeness of the enriched diagrams, which in turn would facilitate automated implementations for tasks such as simulation of noisy quantum systems, fine-tuned quantum circuit optimization techniques for specific quantum devices, or comparison of the effectiveness of different Quantum Error Mitigation techniques. An interesting venue would be to use enrichment over $\mathcal{M}$ to reason about quantum circuit pre- and post-processing techniques, such as circuit cutting [41, 39], in which quantum circuits are “split” into linear combinations of smaller ones that are executed separately. We also believe that it could be possible to integrate monads that capture purely quantum behaviours into our construction to intuitively represent within (enriched) ZX-diagrams phenomena such as superposition of execution orders, like what is done in the Many-Worlds calculus [13].
Strongly related to completeness is to have presentations of the diagrams in terms of generators and equations. We have achieved this by hand in section 5 by using that the algebras for the distribution monad can be presented as convex algebras with a family of operations $+_{p}$. The question is then what the analogue of convex monads is when we use algebras presented by Lawvere theories or sketches [20, 33].
@@ Acknowledgements
First, we write down the diagrams for the natural isomorphisms of the pseudo-monoid that is the image of $(\underline{\mathbf{C}},\odot,U,\alpha,\lambda,\rho,\sigma)$ under $\mathbf{PMon}(G)$, following the diagrams of [17, Section 3].
@@
$\mathbf{PMon}(G)$ maps a pseudo-monoid homomorphism $(h,h^{0},h^{2}):(\underline{\mathbf{C}},\odot_{1},U_{1})\to(\underline{\mathbf{D}},\odot_{2},U_{2})$ to $(G(h),G(h^{0}),G(h^{2}):(G\underline{\mathbf{C}},G(\odot_{1}),G(U_{1}))\to(G\underline{\mathbf{D}},G(\odot_{2}),G(U_{2}))$, as seen from the following unit and multiplication diagrams, for $A\in|\underline{\mathbf{C}}|,B\in|\underline{\mathbf{D}}|$:
@@ Appendix 0.B Monoidal structure in $\mathbf{Alg}^{\mathcal{D}}$
The tensor product of algebras of $\mathcal{D}$ represents bi-convex maps following [2], which we spell out in detail in this section. For algebras $a\colon\mathcal{D}(A)\to A$, $b\colon\mathcal{D}(B)\to B$ and $c\colon\mathcal{D}(C)\to C$ we denote by $\mathrm{BiConv}(a,b;c)$ the set
@@ Appendix 0.E Notation
$\otimes_{T}$ Tensor product in $\mathbf{Alg}^{T}$

### 29 Jan 2024 v3

@@ 4 Enriched Monoidal Categories
Lax monoidal strict $2$-functors $(G,G^{0},G^{2})\colon(\mathbf{V},\otimes,I)\to(\mathbf{W},\times,\ast)$ induce $2$-functors $\mathbf{PMon}(G)\colon\mathbf{PMon}(\mathbf{V},\otimes,I)\to\mathbf{PMon}(\mathbf{W},\times,\ast)$ between $2$-categories of pseudo-monoids, lax homomorphisms and $2$-cells that are compatible with the homomorphism structures. If a $G$ has a monoidal left adjoint $F$, then $\mathbf{PMon}(F)$ is left adjoint to $\mathbf{PMon}(G)$. Finally, if the monoidal categories and functors are symmetric, then the adjunction can be improved to one between symmetric pseudo-monoids.
@@
From this theorem and combining the results from Section 3 we can derive the following corollary, which is our main tool for building monoidal diagrams that are enriched with algebraic operations.
@@
A monoidal monad $T$ on $\mathbf{Set}$ with an adjunction between the free $T$-algebra functor and the underlying category functor (see Corollary 1) gives a free-underlying adjunction $(-)_{0}:\mathbf{Alg}^{T}\text{-}\mathbf{Cat}\rightleftarrows\mathbf{Cat}:F_{*}$. This adjunction lifts to an adjunction between the $2$-categories of symmetric monoidal $\mathbf{Alg}^{T}$-enriched and $\mathbf{Set}$-enriched categories $(-)_{0}:\mathbf{Alg}^{T}\text{-}\mathbf{SMCat}\rightleftarrows\mathbf{SMCat}:F_{*}$. More explicitly, given such a monad $T$ on $\mathbf{Set}$ and a SMC $(\mathbf{C},\otimes,I,\alpha,\lambda,\rho,\sigma)$ we can construct the freely $\mathbf{Alg}^{T}$-enriched SMC $(\underline{\mathbf{C}},\underline{\otimes},\underline{I},\underline{\alpha},\underline{\lambda},\underline{\rho},\underline{\sigma})$.
@@ 5.3 Enriching the Categories $\mathbf{Qubit}$ and $\mathbf{CPM(\mathbf{Qubit})}$
We also define the non-freely enriched category $\underline{\mathbf{CPM(\mathbf{Qubit})}}$ so we can interpret probability distributions as CP maps. For this, we define SMC-structure in the non-free $(\mathbf{Alg}^{\mathcal{D}},\otimes^{\mathcal{D}},\mathcal{D}(\ast))$, with a tensor product of algebras defined by the coequalizer (1), with a more detailed description in Appendix 0.B. The category $(\underline{\mathbf{CPM(\mathbf{Qubit})}},\odot,U)$ has the same objects as $\mathbf{CPM(\mathbf{Qubit})}$ and for every pair of objects $A,B$ the hom-object is an algebra $(\mathbf{CPM(\mathbf{Qubit})}(A,B),\alpha)$ with the $\mathcal{D}$-action turning a formal convex sum of linear maps into an actual sum by scalar multiplication and addition. Composition of hom-objects follows from composition in $\mathbf{CPM(\mathbf{Qubit})}$, and for an object $A$ the identity element is $j_{A}:(\ast,\alpha)\to(\operatorname{Id}_{A},\alpha)$. We define now its symmetric monoidal structure following the definition of enriched SMC from the beginning of Section 4. The tensor product $\odot$ on objects is the same as in $\mathbf{CPM(\mathbf{Qubit})}$, and on hom-objects it is the tensor product in $\mathbf{CPM(\mathbf{Qubit})}$ to the underlying sets: $\odot:(\mathbf{CPM(\mathbf{Qubit})}(A,A^{\prime}),\alpha)\otimes^{\mathcal{D}}(\mathbf{CPM(\mathbf{Qubit})}(B,B^{\prime}),\alpha)$to $(\mathbf{CPM(\mathbf{Qubit})}(A\otimes B,A^{\prime}\otimes B^{\prime}),\alpha)$. The unit $U$ is the one in $\mathbf{CPM(\mathbf{Qubit})}$. The associator, unitors, and symmetry all follow from applying the ones in $\mathbf{CPM(\mathbf{Qubit})}$.
@@ 5.4 Enriched ZX-diagrams and Their Interpretation
We need to take extra care when handling scalars inside the brackets. Indeed, what we have inside distribution brackets is a formal convex sum of ZX-diagrams (or, in the general case, string diagrams), meaning that the SMC rewriting axioms apply to each summand independently. Since summands are also juxtaposed, it might seem like this notation allows for the transfer of scalars from one summand to another. The crux is that, since what is enclosed by trapezoids is a formal sum, we cannot drag scalars from one summand to another using those same monoidal category axioms. This means that we can consider the probabilities (and any scalar factor if present, such as the imaginary unit in Figure 5) to be bound to the wires themselves, and only interact with the ZX-diagrams (or generally string diagrams) that belong to that summand. An alternative is to encapsulate each summand in “bubbles” for stronger visual separation [50, 39].
@@ 6 Discussion and Future Work
In this work, we have shown how to construct freely enriched symmetric monoidal categories over the algebras of a monoidal monad on $\mathbf{Set}$ that satisfies $F\dashv\mathbf{Alg}^{T}(I,-)$ for $F$ the free $T$-algebra functor and $I$ the unit of the monoidal structure, which is the case in particular when $T$ is an affine monad. We have then taken this construction and developed a graphical language that captures the additional algebraic structure of the morphisms for the case of the Distribution monad. We then show how we can use this to study classical probabilistic processes in quantum systems, a highly relevant type of operation for near-term quantum applications. In particular, we extend the ZX-calculus to make it a language for reasoning in an enriched version of $\mathbf{CPM(\mathbf{Qubit})}$.
We believe that this work opens several directions for future research. The most evident one is to prove completeness of the enriched diagrams, which in turn would facilitate automated implementations for tasks such as simulation of noisy quantum systems, fine-tuned quantum circuit optimization techniques for specific quantum devices, or comparison of the effectiveness of different Quantum Error Mitigation techniques. An interesting venue would be to use enrichment over $\mathcal{M}$ to reason about quantum circuit pre- and post-processing techniques, such as circuit cutting [44, 42], in which quantum circuits are “split” into linear combinations of smaller ones that are executed separately. We also believe that it could be possible to integrate monads that capture quantum behaviours into our construction to represent in enriched ZX-diagrams phenomena such as superposition of execution orders, like what is done in the Many-Worlds calculus [14].
Strongly related to completeness is to have presentations of the diagrams in terms of generators and equations. We achieved this by hand in Section 5 by using that the algebras for the distribution monad can be presented as convex algebras with a family of operations $+_{p}$. The question is then what the analogue of convex monads is when using algebras presented by Lawvere theories or sketches [21, 36].
@@ Acknowledgements
First, we write down the diagrams for the natural isomorphisms of the pseudo-monoid that is the image of $(\underline{\mathbf{C}},\odot,U,\alpha,\lambda,\rho,\sigma)$ under $\mathbf{PMon}(G)$, following the diagrams of [18, Section 3].
@@
$\mathbf{PMon}(G)$ maps pseudo-monoid homomorphisms $(h,h^{0},h^{2}):(\underline{\mathbf{C}},\odot_{1},U_{1})\to(\underline{\mathbf{D}},\odot_{2},U_{2})$ to
$(G(h),G(h^{0}),G(h^{2})):(G\underline{\mathbf{C}},G(\odot_{1}),G(U_{1}))\to(G\underline{\mathbf{D}},G(\odot_{2}),G(U_{2}))$
as seen from the following unit and multiplication diagrams, for $A\in|\underline{\mathbf{C}}|,B\in|\underline{\mathbf{D}}|$:
@@ Appendix 0.B Monoidal Structure in $\mathbf{Alg}^{\mathcal{D}}$
The tensor product of algebras of $\mathcal{D}$ represents bi-convex maps following [2], which we give in detail in this section. For algebras $a\colon\mathcal{D}(A)\to A$, $b\colon\mathcal{D}(B)\to B$ and $c\colon\mathcal{D}(C)\to C$ we denote by $\mathrm{BiConv}(a,b;c)$ the set
@@ Appendix 0.E Notation
$\otimes^{T}$ Tensor product in $\mathbf{Alg}^{T}$